1. Electromagnetic radiation travels through vacuum at a speed of ________ m/s.
2. A local radio station often broadcast on the 6-meter band. The frequency of this
electromagnetic radiation is ________ MHz.
3. What is the frequency of light (s-1) that has a wavelength of 3.25 × 10-6
cm?
4. What is the frequency (s-1) of a photon that has an energy of 4.15 × 10-18 J?
5. What is the wavelength ( nm) of a photon that has energy of 4.15 × 10-18 J?
1
Expert's answer
2015-10-26T05:14:48-0400
1. Electromagnetic radiation travels through vacuum at a speed of 2.998x108 m/s.
2. The 6-meter wavelengh corresponds to ν = c/λ = 2.998×108/6 = 5x107 s-1 or 50 MHz
3. The light that has a wavelength of λ = 3.25×10-6 cm corresponds to
ν = c/λ =2.998×108/3.25×10-8 = 9.224×1015 s-1
4. The frequency of a photon that has an energy of 4.15 × 10-18 J is
ν = E/h = 4.15×10-18/(6.63 × 10-34) = 6.26 × 1015 s-1
5. The wavelength of a photon that has energy of 4.15 × 10-18 J is
λ = hc/E = (6.63×10-34 x 2.998×108)/(4.15×10-18) = 4.793 × 10-8 m or 47.93 nm